# Dynamical breaking of inversion symmetry, strong second harmonic generation, and ferroelectricity with nonlinear phonons

### v2

@@ Abstract
We show how crystalline inversion symmetry can be dynamically broken by optical phonons with generic, hardening Kerr-like non-linearities. The symmetry-broken state is reached through a parametric instability that can be accessed by driving close to half the phonon resonance. After the onset of the instability, the system settles to a steady state with inversion-symmetry breaking phonon trajectories and strong second harmonic generation. The time averaged positions of the atoms are displaced relative to equilibrium, resulting in a ferroelectric rectification of the driving signal.
@@ Introduction
The non-equilibrium behavior of condensed matter systems currently attracts considerable attention due to its potential for on-demand control over materials (basov2017_on_demand, ; bloch2022strongly_corr_el_photon, ; rudner2020band, ). Out-of-equilibrium phonons are of particular interest, as lattice distortions have an immediate impact on the electronic properties of solids (forst2011nonlinear_phononics, ; mankowsky2016nonlinear_phononics_2, ). Nonlinear phonon resonances can be expoited to create, enhace and manipulate superconductivity (mankowsky2014_enhanced_superconductivity, ; knap2016phonon_superconduct_theory, ; babadi2017phonon_superconduct_theory, ; cavalleri2018photo_induced_supercond, ; liu2020_res_phonon_light_induced_superc, ), magnetism (fechner2018magnetophononics_review, ; afanasiev2021ultrafast_control_magnetic_phonons, ; disa2023_phonon_ferromag_induce, ; luo2025terahertz_control_magno_phononics, ) and other states of matter (nova2019phonon_ferroelectricity, ; ning2023_phonon_induced_hidden_quadrupolar, ; kaplan2025_spatiotemporal_phonons, ; kaplan2025spatiotemporal_phonons_first_principles, ).
Here, we show how nonlinear phonons can be driven in unconventional ways to create symmetry-forbidden electromagnetic responses and lattice deformations that dynamically break the underlying crystal symmetries. We focus on inversion symmetry, which is known to prohibit second harmonic generation (SHG) and rectification (Boyd_nonlinear_optics, ). We demonstrate that, even for a very generic hardening Kerr-like nonlinearity, this rule can be circumvented by a taylored driving protocol, and a novel symmetry-breaking, non-equilibrium state is reached. For chiral phonons carrying finite angular momentum (juraschek2025chiral_phonons_rev, ) [^1] We refer to phonons carrying finite angular momentum as chiral phonons throughout this letter. We note however that the term axial phonons has been proposed for modes that carry finite angular momentum, but do not propagate., we demonstrate how the strong second harmonic generation leads to diverse Lissajous like non-inversion-symmetric phonon trajectories that create structured magnetic fields influencing the dynamics of electrons and spins on a microscopic level (yaniv2025multicolor, ; luo2023large_magnetic_chiral_phonons, ; juraschek2022giant_magn_field_chiral_phonons, ; xiong2022effective_magn_field_chiral, ). Furthermore a dipolar displacement of atoms from their equilibrium positions in the steady state leads to rectification and ferroelectric response.
A model for nonlinear chiral phonons
We will discuss both, degenerate chiral optical phonons, for which the resonance frequency does not depend on the sense of motion (kahana2024_chiral_rectification, ; cheng2020_dirac_semi_phonon_zeeman, ; mustafa2025origin_phonon_zeeman_mos2, ), as well as phonons with split frequencies for right- and left-handed motion (zhang2015chiral_phonons_hexagonal, ; zhu2018observation_chiral_phonons, ; ishito2023truly_chiral_phonons, ; ueda2023chiral_phonons_xrays, ). We begin with the former case. A simple model for nonlinear, degenerate chiral phonons is given by the Hamiltonian (kahana2024_chiral_rectification, )
$H=P_{x}^{2}+P_{y}^{2}+\frac{\Omega_{0}^{2}}{2}\left(Q_{x}^{2}+Q_{y}^{2}\right)+\frac{\beta}{4}\left(Q_{x}^{2}+Q_{y}^{2}\right)^{2}+V_{\mathrm{l-m}}.$ (1)
@@
where the electric dipole moments of the phonon components are given by $\mathbf{p}_{n}=\mathbf{Z}_{n}Q_{n}$, with effective electric charges $\mathbf{Z}_{n}$. For simplicity, we assume $\mathbf{Z}_{n}\propto\hat{\mathbf{e}}_{n}$. Then, for circularly polarized light, the phonons’ equations of motion read
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{x}E_{x}\cos\omega t$ (3)
$\ddot{Q}_{y}+2\gamma\dot{Q}_{y}+\Omega_{0}^{2}Q_{y}+\beta Q_{y}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{y}E_{y}\sin\omega t,$ (4)
@@
Instability for a single phonon component
To show how the symmetry breaking instability emerges, we first focus on a single phonon component $Q_{x}\left(t\right)$ and set $E_{y}=0$. It is useful to divide $Q_{x}\left(t\right)$ into parts composed of odd and even harmonics:
$Q_{i}\left(t\right)=Q_{i,\mathrm{odd}}\left(t\right)+Q_{i,\mathrm{even}}\left(t\right),$ (5)
@@
$Q_{i,\mathrm{odd}}\left(t+\frac{\pi}{\omega}\right)$ $=-Q_{i,\mathrm{odd}}\left(t\right),$
$Q_{i,\mathrm{even}}\left(t+\frac{\pi}{\omega}\right)$ $=Q_{i,\mathrm{even}}\left(t\right).$ (6)
Naively, one expects $Q_{i,\mathrm{even}}$ to vanish, because the odd-order nonlinearity in the equations of motion (4) does not couple even harmonics to the driving. However, in the following, we describe a route to create even harmonics via a parametric instability.
Let us first study the onset of this instability. Since we assume, for the moment, that $E_{y}=0$, we can set $Q_{y}=0$ and consider the dynamics for the $Q_{x}$ mode alone. Using the decomposition of Eq. (5), we can separate the equation (3) into equations for $Q_{x,\mathrm{odd}}$ and $Q_{x,\mathrm{even}}$. We use that, at the onset of the instability, $Q_{i,\mathrm{even}}$ will be very small, such that $\left|Q_{i,\mathrm{even}}\right|\ll\left|Q_{i,\mathrm{odd}}\right|$. Then, the equation for the odd part, neglecting contributions stemming from $Q_{i,\mathrm{even}}$, reads
@@
This is the equation of a simple driven Duffing oscillator. For our purposes it is sufficient to approximate the response $Q_{x,\mathrm{odd}}$ with the fundamental harmonic and write
@@
$+\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\left[1+h\left(E_{x}\right)\cos\left(2\omega t+2\phi_{x}\right)\right]Q_{x,\mathrm{even}}=0$ (10)
@@
$\tilde{\Omega}_{0}\left(E_{x}\right)=\Omega_{0}\sqrt{1+\frac{3\alpha}{2\Omega_{0}^{2}}F_{x,1}^{2}\left(E_{x}\right)},$ (11)
and $h\left(E_{x}\right)=3\alpha F_{x,1}^{2}\left(E_{x}\right)/\left[2\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\right]$. We used Eq. (8) to approximate $Q_{x,\mathrm{odd}}^{2}$. It is then the constant-in-time part of $Q_{x,\mathrm{odd}}^{2}$ that modifies the resonance frequency of the mode and leads to a blue shift, while the oscillating part of $Q_{x,\mathrm{odd}}^{2}$ acts as a parametric driving for $Q_{x,\mathrm{even}}$. The Mathieu equation (14) is known to exhibit parametric instabilities for $\tilde{\Omega}_{0}\left(E_{x}\right)=n\omega$, with $n$ a positive integer (Landau_Lifshitz_Mechanics, ). However, $Q_{x,\mathrm{even}}$has to obey Eq. (6), which excludes the $n=1$ resonance. The $n=2$ resonance, however, is allowed, and leads to the symmetry-breaking instability we want to study. Here $\tilde{\Omega}_{0}\left(E_{x}\right)=2\omega$, such that for driving slightly above half the original resonance frequency of $\Omega_{0}$, we expect a response at $\tilde{\Omega}_{0}\left(E_{x}\right)$ – i.e., we expect strong SHG.
As is typical for parametric resonances, the instability occurs in a small frequency window where for $\Delta=2\omega-\tilde{\Omega}_{0}$ holds (see e.g. (turyn1993damped_Mathieu, ), p. 394  [^2] The equations on p. 390 of this reference contain a typo.)
@@
Here, and in what follows, we omit writing out the $E_{x}$-dependence of $F_{x,1}$, $\tilde{\Omega}_{0}$, $h$ and $\Delta$ explicitly, except when it is needed for clarity. The blue shift and the instability window are illustrated with the results of a numerical simulation in Fig. 1 a).
To overcome damping effects, a minimal driving amplitude is required. This threshold amplitude $E_{x,*}$ can be calculated by setting $\Delta=0$. To leading order in $\gamma/\Omega$, we find $h\left(E_{x,*}\right)=\sqrt{8\gamma/\Omega_{0}}$ where $\Omega_{0}$ is the $E_{x}$-independent resonance frequency of Eq. (1). This expression can be inverted for $E_{x}$ using Eq. (9). For small damping, the threshold electric field amplitude is then given by
$E_{x,*}\approx\frac{\sqrt{3}\Omega^{3}}{2^{3/4}\sqrt{\beta}Z_{x}}\left(\frac{\gamma}{\Omega}\right)^{1/4}$ (13)
As can be expected, $E_{x,*}$ is lowered by a strong nonlinearity $\beta$ and increased by a larger $\gamma$. The result of Eq. (13) is confirmed by numerical simulations as shown in Fig. 1 c).
Upon going through the parametric instability at $2\omega\approx\tilde{\Omega}_{0}\left(E_{x}\right)$, the phonons reach a stable trajectory. This behavior is not uncommon in nonlinear systems (dykman1998fluctuational_transitions_parametric_osc, ; marthaler2006parametric_osc_quantum_switching, ), however, in our case the steady state is characterized by strong fundamental and second harmonic response, as well as a considerable DC offset. The spectrum of $Q_{x}$ in this steady state, obtained by solving Eq. (3) numerically with $E_{y}=0$, is shown in Fig. 1 b). We note in passing, that the instability and steady state studied here are known in nonlinear systems literature, although the only extensive study, to our knowledge, is presented in Ref. (ys1991_duffing_symm_breaking_SHG, ). Before extending our results to the chiral system of Eqs. (3) and (4), we investigate the DC component of the steady-state leading to a ferroelectric response, and show how auxiliary phonon modes can be exploited to trigger the symmetry breaking instability resonantly.
@@
Ferroelectricity
We now show that the instability and steady state outlined above, necessarily imply the presence of a static displacement of the atoms from their equilibrium positions. To see this, we averagy Eq. (3) over one period of the drive. As above, we assume $E_{y}=0$ and therefore $Q_{y}=0$. Writing $Q_{x}\left(t\right)=\sum_{a}F_{x,a}\cos\left(n\omega t+\varphi_{x,a}\right)$, and truncating the series at $n=2$, find
@@
$\quad+\left.\frac{3}{4}F_{x,1}^{2}F_{x,2}\cos\left(2\varphi_{x,1}-\varphi_{x,2}\right)+F_{x,0}^{3}\right].$
@@
$F_{x,0}=-\frac{3\beta}{4\omega_{0}^{2}}F_{x,1}^{2}F_{x,2},$
@@
showing that any response at the second harmonic is accompanied by a DC offset. Being third order in the first and second harmonic amplitudes, we expect the DC offset to be smaller in magnitude, it can however, still be sizable [see Fig. (1) b)]. We conclude that although the symmetry breaking instability is triggered by an oscillating driving field, inversion symmetry is still statically broken. This results in constant in time electric fields produced by the dipoles $\mathbf{p}=Z_{x}F_{x,0}$, where $Z_{x}$ is the effective electric charge of the phonon mode in question. The driving signal is thus rectified and triggers a ferroelectric response.
Exploiting resonant modes

### v3

@@ Abstract
We show how crystalline inversion symmetry can be dynamically broken by optical phonons with generic, hardening Kerr-like non-linearities. The symmetry-broken state is reached through a parametric instability that can be accessed by driving close to half the phonon frequency. The system then settles to a steady state with inversion-symmetry breaking phonon trajectories and strong second harmonic generation. The time averaged positions of the atoms are displaced relative to equilibrium, resulting in a ferroelectric rectification of the driving signal. For circularly polarized phonons, complex Lissajous-like trajectories can be achieved.
@@ Introduction
The non-equilibrium behavior of condensed matter systems currently attracts considerable attention due to its potential for on-demand control over materials (basov2017_on_demand, ; bloch2022strongly_corr_el_photon, ; rudner2020band, ). Out-of-equilibrium phonons are of particular interest, as lattice distortions have an immediate impact on the electronic properties of solids (forst2011nonlinear_phononics, ; mankowsky2016nonlinear_phononics_2, ). Nonlinear lattice dynamics can be expoited to enhace and manipulate superconductivity (mankowsky2014_enhanced_superconductivity, ; knap2016phonon_superconduct_theory, ; babadi2017phonon_superconduct_theory, ; cavalleri2018photo_induced_supercond, ; liu2020_res_phonon_light_induced_superc, ), magnetism (fechner2018magnetophononics_review, ; afanasiev2021ultrafast_control_magnetic_phonons, ; disa2023_phonon_ferromag_induce, ; luo2025terahertz_control_magno_phononics, ) and other states of matter (nova2019phonon_ferroelectricity, ; ning2023_phonon_induced_hidden_quadrupolar, ; kaplan2025_spatiotemporal_phonons, ; kaplan2025spatiotemporal_phonons_first_principles, ). Other effects predicted for driven, nonlinear collective modes include non-trivial steady states (kiselev2024MFPD_plasmons_nat, ; kiselev2024_PRB_MFPD, ; kiselev2025exciting, ; wanic2025entanglement, ; kaplan2025_spatiotemporal_phonons, ; kaplan2025spatiotemporal_phonons_first_principles, ; kaplan2025optically_induced_goldstone_faraday, ), chaos (wanic2024magnetoelectric, ), and parametric amplification (cartella2018parametric, ; buzzi2021higgs_lasing, ; michael2024photonic, ).
Here, we show how IR-active phonons oscillating in anharmonic interatomic potentials can be driven in unconventional ways to break underlying crystal symmetries and create symmetry-forbidden electromagnetic responses. We focus on inversion symmetry, which is known to prohibit second harmonic generation (SHG) and rectification (Boyd_nonlinear_optics, ). We demonstrate that, for a very generic hardening Kerr-like nonlinearity, this rule can be circumvented by a taylored driving protocol, resulting in a symmetry-breaking steady state with strong phononic SHG and DC responses. We stress that no double well-structure with metastable ferroelectric states, as e.g. exploited in Ref. (li2019_driven_metastable_ferroelectricity, ), is required.
The DC response leads to a static displacement of atoms from their equilibrium positions, generating a constant-in-time dipolar electric field. This behavior is analogous to that of an equilibrium ferroelectric, where – below a critical temperature – dipolar fields form due to spontaneous, symmetry-breaking lattice deformations. In our case, the static dipolar response is triggered by a dynamical breaking of inversion symmetry, and we argue that the predicted effect can be used to create an on-demand ferroelectric. In particular, we show that the non-equilibrium ferroelectric response exhibits hysteretic behavior, which underlines that the effect is not merely a nonlinear piezoelectric response (disa2020polarizing_symmetry_breaking_light_nonlinearity, ), but instead a symmetry-breaking out-of-equilibrium phase. We show that this phase is robust to noise, and is accessible at reasonable field strenghts.
We then explore the above effects in a system of circularly polarized phonons (juraschek2025chiral_phonons_rev, ), which are sometimes referred to as chiral phonons (kahana2024_chiral_rectification, ), and have been used to manipulate magnetic states. Here, the strong SHG leads to Lissajous-like, non-inversion-symmetric phonon trajectories. Such trajectories create structured magnetic fields influencing the dynamics of electrons and spins on a microscopic level (luo2023large_magnetic_chiral_phonons, ; juraschek2022giant_magn_field_chiral_phonons, ; xiong2022effective_magn_field_chiral, ), and provide an interesting way of engineering electronic behavior (yaniv2025multicolor, ).
Symmetry-breaking instability and second harmonic generation
Figure 1: a) Phonons [Eq. (18)] driven with a linearly polarized electric field $E_{x}\cos\left(\omega t\right)$ oscillating at frequency $\omega=0.6\Omega_{0}$. The effective phonon frequency $\tilde{\Omega}_{0}\left(E_{x}\right)$, given in Eq. (10), exhibits a blue shift as $E_{x}$ is increased. Around $\tilde{\Omega}_{0}\left(E_{x}\right)\approx 2\omega$, the system enters the symmetry breaking state with strong second harmonic generation. In the symmetry-breaking regime, the resonance curve $\tilde{\Omega}_{0}\left(E_{x}\right)$ is interrupted. A small damping of $\gamma=0.001$ was chosen to have a large instability interval for clarity. b) Spectrum of $Q_{x}\left(t\right)$ in the symmetry breaking state. c) Amplitude ratio of second and first harmonics across the symmetry breaking transition for different dampings $\gamma$. d) Hysteresis of the mean polarization $p_{x}=Z_{x}F_{x,\text{0}}=\left(2\pi/\omega\right)\int_{0}^{2\pi/\omega}dt\,Q_{x}\left(t\right)$ when a static bias field $E_{\mathrm{\mathrm{DC}}}$ is swept through positive and negative values in a triangle-wave-like manner.
To show how the symmetry breaking instability emerges, we first focus on a single IR-active phonon mode $Q_{x}\left(t\right)$. For a generic, inversion symmetric system, the dynamics of $Q_{x}\left(t\right)$ is governed by the Hamiltonian
$H=\frac{1}{2}P_{x}^{2}+\frac{\Omega_{0}^{2}}{2}Q_{x}^{2}+\frac{\beta}{4}Q_{x}^{4}+V_{\mathrm{l-m}}.$ (1)
@@
$V_{\mathrm{l-m}}=-\mathbf{E}\cdot\mathbf{p}_{x},$ (2)
where the electric dipole moment of the phonon is given by $\mathbf{p}_{x}=\mathbf{Z}_{x}Q_{x}$, with effective charge $\mathbf{Z}_{x}$. We assume $\mathbf{Z}_{x}\propto\hat{\mathbf{e}}_{x}$. Then, for circularly polarized light, the phonon equation of motion reads
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}^{3}$ $=Z_{x}E_{x}\cos\omega t$ (3)
@@
It is often assumed that the third order nonlinearity of Eq. (3) does not generate even harmonics of the driving frequency $\omega$. Indeed, shifting the time coordinate according to $t\rightarrow t+\frac{\pi}{\omega}$ changes the sign of the right hand side, which seems to enforce the symmetry $Q_{x}\left(t+\frac{\pi}{\omega}\right)=-Q_{x}\left(t\right)$, excluding any response at frequencies $2n\omega$, where $n\in\mathbb{N}$, including zero. We now show, that this symmetry, stemming from the inversion symmetry of the Hamiltonian (1), can be dynamically broken.
It is useful to divide $Q_{x}\left(t\right)$ into parts composed of odd and even harmonics:
$Q_{x}\left(t\right)=Q_{x,\mathrm{odd}}\left(t\right)+Q_{x,\mathrm{even}}\left(t\right),$ (4)
@@
$Q_{x,\mathrm{odd}}\left(t+\frac{\pi}{\omega}\right)$ $=-Q_{x,\mathrm{odd}}\left(t\right),$
$Q_{x,\mathrm{even}}\left(t+\frac{\pi}{\omega}\right)$ $=Q_{x,\mathrm{even}}\left(t\right).$ (5)
We now show how even harmonics ($Q_{i,\mathrm{even}}$) can be created via a parametric instability. Let us first study the onset of this instability. Using the decomposition of Eq. (4), we can separate Eq. (A 7) into cupled equations for $Q_{x,\mathrm{odd}}$ and $Q_{x,\mathrm{even}}$. We use that, at the onset of the instability, $Q_{i,\mathrm{even}}$ will be very small, such that $\left|Q_{i,\mathrm{even}}\right|\ll\left|Q_{i,\mathrm{odd}}\right|$. Then, the equation for the odd part, neglecting contributions stemming from $Q_{i,\mathrm{even}}$, reads
@@
For our purposes it is sufficient to approximate the response $Q_{x,\mathrm{odd}}$ with the fundamental harmonic and write
@@
$+\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\left[1+h\left(E_{x}\right)\cos\left(2\omega t+2\varphi_{x}\right)\right]Q_{x,\mathrm{even}}=0$ (9)
@@
$\tilde{\Omega}_{0}\left(E_{x}\right)=\Omega_{0}\sqrt{1+\frac{3\beta}{2\Omega_{0}^{2}}F_{x,1}^{2}\left(E_{x}\right)},$ (10)
and $h\left(E_{x}\right)=3\alpha F_{x,1}^{2}\left(E_{x}\right)/\left[2\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\right]$. We used Eq. (7) to approximate $Q_{x,\mathrm{odd}}^{2}$. It is then the constant-in-time part of $Q_{x,\mathrm{odd}}^{2}$ that modifies the resonance frequency of the mode and leads to a blue shift, while the oscillating part of $Q_{x,\mathrm{odd}}^{2}$ acts as a parametric driving for $Q_{x,\mathrm{even}}$. The Mathieu equation (A 8) is known to exhibit parametric instabilities for $\tilde{\Omega}_{0}\left(E_{x}\right)=n\omega$, with $n$ a positive integer (Landau_Lifshitz_Mechanics, ). However, $Q_{x,\mathrm{even}}$has to obey Eq. (5), which excludes the $n=1$ resonance. The $n=2$ resonance, however, is allowed, and leads to the symmetry-breaking instability we want to study. Here $\tilde{\Omega}_{0}\left(E_{x}\right)=2\omega$, such that for driving slightly above half the original resonance frequency of $\Omega_{0}$, we expect a response at $\tilde{\Omega}_{0}\left(E_{x}\right)$ – i.e., we expect strong SHG.
As is typical for parametric resonances, the instability occurs in a small frequency window where for $\Delta=2\omega-\tilde{\Omega}_{0}$ holds (see e.g. (turyn1993damped_Mathieu, ), p. 394 (Note2, ))
@@
Here, and in what follows, we omit writing out the $E_{x}$-dependence of $F_{x,1}$, $\tilde{\Omega}_{0}$, $h$ and $\Delta$ explicitly, except when it is needed for clarity. The blue shift and the instability window are shown with the results of a numerical simulation in Fig. 1 a) (see AppendixAppendix.1.1 for a note on the initial conditions used). We stress that, while in our analytical calculation we approximated the odd response by a single harmonic and linearized in $Q_{x,\mathrm{even}}$, we used the full Eq. (3) for the simulation. As can be expected there are discrepancies due to the approximations made, however, the instability window is clearly identifiable.
To overcome damping effects, a minimal driving amplitude is required. This threshold amplitude $E_{x,*}$ can be calculated by setting $\Delta=0$. To leading order in $\gamma/\Omega$, we find $h\left(E_{x,*}\right)=\sqrt{8\gamma/\Omega_{0}}$. This expression can be inverted for $E_{x}$ using Eq. (8). For small damping, the threshold electric field amplitude is then given by
$E_{x,*}\approx\frac{\sqrt{3}\Omega^{3}}{2^{3/4}\sqrt{\beta}Z_{x}}\left(\frac{\gamma}{\Omega_{0}}\right)^{1/4}.$ (12)
Notice that $\Omega_{0}$ is the bare resonance frequency, but the blue-shift has been taken into account in the derivation of Eq. (12). As can be expected, $E_{x,*}$ is lowered by a strong nonlinearity $\beta$ and increased by a larger $\gamma$. The result of Eq. (12) is confirmed by numerical simulations as shown in Fig. 1 c). The threshold fields for different (bare) phonon resonance frequencies $\Omega_{0}$ are given in Table 1. Since $E_{x,*}\sim\Omega_{0}^{3}$ holds, much stronger amplitudes are required at higher frequencies. The amplitudes are, however, in an experimentally accessible range.
$\Omega_{0}/2\pi$ $1\,\mathrm{THz}$ $3\,\mathrm{THz}$ $5\,\mathrm{THz}$
$E_{x,*}$ $1.5\cdot 10^{6}\,\mathrm{V/m}$ $4.1\cdot 10^{7}\,\mathrm{V/m}$ $1.9\cdot 10^{8}\,\mathrm{V/m}$
@@
Symmetry breaking steady-state and pulsed excitation
The above analysis demonstrates that the usual solution, consisting of a strong response at the fundamental harmonic and perturbative corrections at odd harmonics, is unstable if driving frequency and strength are tuned such that the condition of Eq. (11) at $2\omega\approx\tilde{\Omega}_{0}\left(E_{x}\right)$ is fulfilled. This instability is associated with strong second harmonic generation due to the parametric resonance in Eq. (A 8). Investigating the dynamics of the nonlinear phonon mode in this regime, we find that the system converges to a stable steady-state which breaks the inversion symmetry of the system. Besides a strong second harmonic, this symmetry-breaking is characterized by a relatively large DC offset. The spectrum of $Q_{x}$ in the steady state, obtained by solving Eq. (3) numerically, is shown in Fig. 1 b). We note in passing, that the instability and steady state studied here have been described in nonlinear systems literature, although the only extensive study, to our knowledge, is presented in Ref. (ys1991_duffing_symm_breaking_SHG, ).
For low frequency phonons the instability is accessible already at relatively small driving fields (see Table1), which might be produced by continuous wave lasers. However, reaching the symmetry breaking with pulsed signals is also possible. We performed a series of numerical experiments using the same parameters as in the continuous wave case, and find that short pulses with as few as five oscillation cycles (see Appendix Appendix.1) are sufficient to observe both second harmonic generation and a DC offset. However, flat-top pulses must be used, because, due to the nonlinear blue-shift, the frequency and amplitude of the drive must be tuned toghether. The dashed lines in Fig.1 b) show the response spectrum obtained for a flat top pulse of five optical cycles. The finite duration of the pulse introduces spectral broadening, however, the SHG and the DC offset are clearly visible.
Before extending our results to a system of chiral phonons, we investigate how the DC component of the steady-state at $2\omega\approx\tilde{\Omega}_{0}\left(E_{x}\right)$ leads to a ferroelectric response, and show how auxiliary phonon modes can be exploited to trigger the symmetry breaking instability resonantly.
Ferroelectricity, hysteresis, stability to noise
We now show that the symmetry-breaking steady-state outlined above, necessarily implies the presence of a static displacement of the atoms from their equilibrium positions. To see this, we average Eq. (3) over one period of the drive. Writing $Q_{x}\left(t\right)=\sum_{a}F_{x,a}\cos\left(a\omega t+\varphi_{x,a}\right)$, and truncating the series at $a=2$, we find
@@
$\quad+\left.\frac{3}{4}F_{x,1}^{2}F_{x,2}\cos\left(2\varphi_{x,1}-\varphi_{x,2}\right)+F_{x,0}^{3}\right].$ (13)
@@
$F_{x,0}=-\frac{3\beta}{4\omega_{0}^{2}}F_{x,1}^{2}F_{x,2},$ (14)
showing that any response at the second harmonic is accompanied by a DC offset. Being third order in the first and second harmonic amplitudes, we expect the DC offset to be smaller in magnitude, it can however, still be sizable [see Fig. (1) b)].
@@
We conclude that although the symmetry breaking instability is triggered by an oscillating driving field, inversion symmetry is statically broken. This results in constant-in-time electric fields produced by the dipoles $\mathbf{p}=\mathbf{Z}_{x}F_{x,0}$, where $\mathbf{Z}_{x}$ is the effective electric charge of the phonon mode in question. The driving signal is thus rectified and triggers a ferroelectric response.
This dynamically triggered ferroelectric response shares significant similarities with equilibrium ferroelectrics, showing that it can be used as a road to on-demand light driven ferroelectricity. To substantiate this claim, we demonstrate that hysteresis can be observed when the phonon mode is in the symmetry-breaking state. To this end we add a static component $E_{\mathrm{DC}}$ to the driving field in Eq. (3) and sweep it through positive and negative values in a triangle-wave-like manner. We numerically integrate Eq. (3) while slowly changing $E_{\mathrm{DC}}$. Fig.1d) shows the mean polarization
$p_{x}=Z_{x}F_{x,\text{0}}=\left(2\pi/\omega\right)\int_{0}^{2\pi/\omega}dt\,Q_{x}\left(t\right),$ (15)
which exhibits a hysteresis. Interestingly, the hysteresis curve is pinched – also a feature known from equilibrium ferroelectrics (xu2016pinched_hysteresis_ferroelectrics, ).
[…]
Figure 2: Collapse of the symmetry-breaking steady state under the influence of noise. In the case of thermal noise and $\Omega_{0}/2\pi=3\,\mathrm{THz}$, the critical noise strength $A\approx 0.1Z_{x}E_{x}$ corresponds to a temperature of $\sim 700\,\mathrm{K}$. The same parameters as in Fig.1 were used.
Exploiting auxiliary IR- and Raman-active modes